40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7837     Entry time: Mon Dec 17 11:20:58 2012
Author: Jenne 
Type: Update 
Category: SUS 
Subject: Beam dumps on vertex oplevs removed 

I'm not sure when this was done, but there were beam dumps in front of the lasers for BS/PRM oplevs as well as ITMY/SRM oplevs.  MICH wasn't holding lock very nicely, so I poked around, and the Sum values for all of these optics' oplevs seemed too low, so I went to look, and found dumps.  I have removed these, and now BS and ITMY oplevs are back to normal.  (PRM and SRM are still misaligned right now, so I'll check those later, but they should be fine).

BS's oplev has been enabled while non-existant, at least for the whole weekend, since I found it enabled.  ITMY I found misaligned, so it's oplev servos were off.

In other news, we should get back in the habit of restoring all optics before we leave for the night / whenever locking activities are finished. 

ELOG V3.1.3-